solve each system of equations algebraically
5x-2y=20
2x+3y=27

Answer by user_dude2008(1862)

5x-2y=20
2x+3y=27


10x-4y=40
-10x-15y=-135

-19y=-95
y=5

2x+3(5)=27
2x+15=27
2x=12
x=6
Answers: x=6, y=5